# The Preheating Stage on The Starobinsky Inflation after ACT

### v1

@@ 8.4 Reheating temperature by preheating
As we learn from the perturbative reheating, it has some issues at certain points. Thus, the reheating temperature should be obtained from the burst of particle production due to resonance. However, before we proceed, we need to assume the most possible mechanism.
Our chosen mechanism is as follows: Due to the nature of the matter-dominated regime, it is expected that the reheating temperature will be evaluated at the end of the preheating stage. However, the reheating temperature should not be evaluated from the ordinary perturbative decay of either bosons or fermions, as discussed in the previous subsection. Instead, the reheating temperature should be evaluated from the last burst of scalaron oscillation that produces scalar $\chi$ at the end of the preheating stage [^9] This was discussed in our previous paper (see Ref. [30]), although we used the beginning of the second stage of preheating.. The produced scalar $\chi$ will immediately decay into relativistic fermions $\psi$ to reheat the universe [^10] We choose the bosons $\chi$ as the product at the zero crossing of the scalaron due to the abundance of bosons over fermions during preheating. Fermions $\psi$ are chosen as the decay product of $\chi$ because they are more directly related to SM particles due to their chiral nature.. Intuitively, we can assume that during this time, the remaining energy density of the inflaton is very small. Thus, a low reheating temperature is expected.
@@
where $\delta\rho$ is the energy density produced by the single crossing during the ’last’ preheating. The $\delta\rho$ can be evaluated as [30, 31]
[…]
@@
Considering the chosen parameters, we need to clarify some favored values. The coupling $y$ should be small to suppress instant reheating (see [19]) [^11] Fortunately, if we set $y$ to be small, our coupled e.o.m. (3.2) and (3.3) do not necessarily consider the growth of $\psi$.. For $m_{\chi}=10^{-15}M_{p}\sim 10^{-16}M_{p}$, this lies close to the electroweak scale. A model with this mass range favors linkage to the SM. Interestingly, a mass of this value does not contradict our model. In addition, since $m_{\psi}\ll m_{\chi}$ is required to allow for smooth decay, it is clear that $m_{\psi}$ is on the order of SM particle masses. For our chosen parameters, $\chi$ could correspond to the Higgs boson, and $\psi$ could be SM fermions [^12] The small Yukawa coupling may naturally disfavor some large quark couplings.. However, we must be aware that due to the large couplings of the Higgs to gauge bosons (and some large-coupling fermions), such interactions could spoil our low-reheating inflationary model. Finally, it should be noted that both $\chi$ and $\psi$ masses should lie above the electroweak scale, as neither fields do not behave like an SM particle. Note that $y$ could be much smaller. Thus, all other parameters could be set to higher values, $k_{\chi}\gg 10^{-8}M_{p}$ and $m_{\chi}\gg 10^{-16}M_{p}$.
@@ 9 Conclusion
In this paper, we reinvestigate Starobinsky inflation, particularly the preheating stage, following the release of the ACT results. These results suggest a significant increase in the e-fold number ($N\simeq 75$) and constrain the reheating temperature to $T_{\text{reh}}\sim 10^{4}$ GeV if we adopt the $68\%$ confidence level. Consequently, the preheating and reheating mechanisms must be modified to align with these new results. Such a low reheating temperature affects the entire preheating stage, not just the reheating mechanism. In addition, the direct perturbative decay from the scalaron to scalar $\chi$ is strongly disfavored if the preheating mechanism is dominant and the high reheating temperature is implied. On the other hand, reheating via the decay of the scalaron to fermions $\psi$ is also disfavored if the preheating stage is dominant, although less severely, due to the possibility of achieving a lower reheating temperature and its chirality nature.
@@
In the preheating stage, we consider two mechanisms, which we refer to as Case 1 and Case 2, distinguished by the presence of tachyonic conditions. Case 1 describes the preheating stage when the tachyonic condition is not applied. In this mode, preheating is shown to be inefficient. Consequently, the preheating stage lasts much longer than the analytical prediction ($\gg 10^{5}\pi/M$). On the contrary, Case 2 exhibits purely tachyonic conditions. In this mode, the preheating stage ceases rapidly due to efficient particle production ($\ll 10^{5}\pi/M$). Neither the numerical prediction from Case 1 nor that from Case 2 alone satisfies the analytical calculation. For this reason, the realistic preheating mechanism may involve a combination of both Case 1 and Case 2. To accommodate this idea, at least one momentum, $k_{\phi}$ or $k_{\chi}$, should not be fixed but instead follow a time-dependent function that disrupts the tachyonic conditions. This is referred to as the ceased-out mechanism [30].
The disruption of the tachyonic conditions can occur through two mechanisms, based on the following question: At which stage does tachyonic preheating occur—during the early or late phase of preheating? This question arises particularly when we consider PBH production during preheating. It has been shown that the appropriate PBH production occurs during the early stage of preheating. Based on this idea, it is clear that a preheating stage where tachyonic condition intervenes on the late stage is strongly disfavored.
We have also considered the favored parameters in our model based on our discussion in the previous sections. In our model, $y$ should be small to avoid instant reheating, which would lead to a high reheating temperature. The non-minimal coupling $\xi$ should be of order unity or much lower due to unitarity issues (in this paper, we set $\xi=1$). The mass of $\chi$ should be lower than $\frac{1}{2}M$ to ensure that the scalaron decays completely. In addition, $m_{\chi}$ should be higher than the electroweak scale ($\gtrsim 10^{3}$ GeV $\approx 10^{-15}M_{p}$), as $\chi$ is shown to be different to the Higgs field to preserve the viability of our model [^13] If $m_{\chi}$ is at the electroweak scale and not the Higgs field, it may have already been observed.. The mass $m_{\psi}$ should be much lower than $m_{\chi}$ to ensure a successful reheating scenario [^14] As $\psi$ is supposed to reheat the universe, it should be relativistic. Thus, $m_{\psi}\ll m_{\chi}$ is preferred.. The parameters $k_{\phi}$ and $k_{\chi}$ in the early stage of preheating should satisfy the tachyonic condition (see eq. (3.12)). One should note that, as $\chi$ in the early stage of preheating is heavy non-relativistic, it is implied that $k_{\chi}$ should be low. On the contrary, in the late stage, they should violate the tachyonic condition. Finally, the evolution of $k_{\phi}$ is favorable for PBH production. One can see table 3 for the appropriate values of the parameters. Note that all parameters in 3 are under the constraint of the reheating temperature.
@@
$m_{\chi}$ $10^{-6}\sim 10^{-15}M_{p}$ $<\frac{1}{2}M$
$m_{\psi}$ $10^{-7}\sim 10^{-15}M_{p}$ $\ll m_{\chi}$
$k_{\phi}k_{\chi}$ $>8.23\times 10^{-11}M_{p}^{2}$ $>\xi M^{2}$, early preheating
$k_{\phi}k_{\chi}$ $<8.23\times 10^{-11}M_{p}^{2}$ $<\xi M^{2}$, end preheating
Table 3: The viable values of the parameters in our model. Note that we set $\xi=1$.

### v3

@@ 7.4 Reheating temperature by preheating
As we learn from the perturbative reheating, it has some issues at certain points. Thus, the reheating temperature should be determined from the burst of particle production resulting from resonance. Our chosen mechanism is as follows: Due to the nature of the matter-dominated regime, it is expected that the reheating temperature will be evaluated at the end of the preheating stage. However, the reheating temperature should not be evaluated from the ordinary perturbative decay of either bosons or fermions, as discussed in the previous subsections. Instead, the reheating temperature should be evaluated from the last burst of scalaron oscillation that produces scalar $\chi$ at the end of the preheating stage [^7] This was discussed in Ref. [risdianto2025second]). Although in this reference, it used the beginning of the second stage of preheating.. The produced scalar $\chi$ will immediately decay into relativistic fermions $\psi$ to reheat the universe [^8] We choose the bosons $\chi$ as the product at the zero crossing of the scalaron due to the abundance of bosons over fermions during preheating. Fermions $\psi$ are chosen as the decay product of $\chi$ because they are more directly related to SM particles due to their chiral nature.. Intuitively, we can assume that during this time, the remaining energy density of the inflaton is very small. Thus, a low reheating temperature is expected.
@@
where $\delta\rho$ is the energy density produced by the single crossing during the ’last’ preheating. The $\delta\rho$ can be evaluated as [risdianto2025second, hashimoto2021inflation]
$\delta\rho=\int^{\pi M^{-1}}_{0}dt\hskip 2.84526pt\Gamma_{\chi\rightarrow\bar{\psi}\psi}n_{\chi}\bar{m}_{\chi}e^{-\int^{t}_{0}\Gamma_{\chi\rightarrow\bar{\psi}\psi}dt^{\prime}},$ (7.10)
@@
Considering the chosen parameters, we need to clarify some favored values. The coupling $y$ should be small to suppress instant reheating (see [bezrukov2009initial]). For $m_{\chi}$, this lies close to the electroweak scale and should be higher. A model with this mass range favors linkage to the SM. However, as $\chi$ and fermion $\psi$ have bare masses, and their mass does not come from the symmetry breaking, they should not belong to the SM particles. [^9] As it is noted that the SM particles are massless at higher than the electroweak scale Thus, for safety, $m_{\chi}\gg m_{\psi}\gg 10^{-16}M_{p}$. In conclusion, both $\chi$ and $\psi$ masses should lie above the electroweak scale, as neither fields behave like an SM particle. Note that $y$ could be much smaller. Thus, all other parameters could be set to higher values, $k_{\chi}\gg 10^{-8}M_{p}$ and $m_{\chi}\gg 10^{-16}M_{p}$. In this model, it seems that the reheating temperature is weakly constrained by parameters $y$, $k_{\chi}$, and $m_{\chi}$.
Lastly, Table 2 presents the preferred parameter values that accommodate the constrained reheating temperature. From the table, it can be seen that a reheating temperature of approximately $10^{4}$ GeV can be achieved. These parameters can also be adjusted to yield a higher reheating temperature (e.g., $10^{8}$ GeV), which is consistent with the previous Planck results. However, when applying the bound $n_{s}=0.9743\pm 0.0034$, the expected reheating temperature decreases to only a few hundred MeV. Therefore, our proposed mechanism fails under this condition and requires further investigation in future work [Risdianto2026PreheatingStarobinsky].
@@ 8 Conclusion
In this paper, we reinvestigate Starobinsky inflation, particularly the preheating stage, following the release of the ACT results. These results suggest a significant increase in the e-fold number ($N\simeq 75$) and constrain the reheating temperature to $T_{\text{reh}}\sim 10^{4}$ GeV if we adopt the center value with $68\%$ of the confidence level. Consequently, the preheating and reheating mechanisms must be modified to align with these new results. Such a low reheating temperature affects the entire preheating stage, not just the reheating mechanism. In addition, the direct perturbative decay from the scalaron to scalar $\chi$ is strongly disfavored if the preheating mechanism is dominant and the high reheating temperature is implied. On the other hand, reheating via the decay of the scalaron to fermions $\psi$ is also disfavored if the preheating stage is dominant, although less severely, due to the possibility of achieving a lower reheating temperature and its chirality nature.
For the efficient preheating to occur, a certain initial condition of $\chi$ needs to exist at the start of preheating. This initial condition is believed to have existed because $\chi$ is a spectator field during inflation. Surprisingly, this burst of resonance particles does not occur during the inflaton zero-crossing, as many inflationary models do, but rather at the moment when tachyonic preheating takes place. However, tachyonic preheating itself does not seem to be efficient enough to drain the inflaton’s energy density at once. This is a rather peculiar condition compared with many other models of inflation.
@@
In the following, we have also considered the favored parameters in our model based on our discussion in the previous sections. In our model, $y$ should be small to avoid instant reheating, which would lead to a high reheating temperature. The non-minimal coupling $\xi$ should be treated carefully, since the preheating could not be efficient enough if $\xi\simeq 1$ and should not be greater than $10$ to avoid the unitarity issues. Also, $m_{\chi}$ should be lower than $\frac{1}{2}M$ to ensure that the scalaron decays completely and higher than the electroweak scale ($\gtrsim 10^{3}$ GeV $\approx 10^{-15}M_{p}$), as $\chi$ is shown to be different to the SM particles to preserve the viability of our model [^10] If $m_{\chi}$ is at the electroweak scale and not the Higgs field, it may have already been observed. Also, the mass is not coming from the symmetry breaking in which should not belong to SM particles in the first place.. The mass $m_{\psi}$ should be much lower than $m_{\chi}$ to ensure a successful reheating scenario [^11] As $\psi$ is supposed to reheat the universe, it should be relativistic. Thus, $m_{\psi}\ll m_{\chi}$ is preferred.. One should note that, as $\chi$ in the early stage of preheating is a heavy, non-relativistic particle, it is implied that $k_{\chi}$ should be much lower than $\xi M^{2}$. Finally, Table 3 provides the appropriate values of the parameters. It is also shown that the chosen parameters could allow both $\chi$ and $\psi$ to be detected in future colliders.
In addition, our preheating mechanism successfully describes scenarios with a lower reheating temperature but fails when the temperature drops below $10^{3}$ GeV. The uncertainty in the ACT result, $n_{s}=0.9743+0.0034$, could lead to a reheating temperature on the order of below the GeV scale (approximately $10^{2}$ MeV), under which our preheating mechanism no longer works. This issue will be addressed in future work.
@@
$\chi(0)$ $\sim 10^{-3}M_{p}$ efficient preheating
$\xi$ $\approx 10$ constraint: unitarity issue (upper bound), and efficient preheating (lower bound)
$m_{\chi}$ $10^{-6}\sim 10^{-14}M_{p}$ $\frac{1}{2}M$ (no inflaton’s remnant condition) and $\gg$ electroweak scale
$m_{\psi}$ $10^{-7}\sim 10^{-15}M_{p}$ $\ll m_{\chi}$ and $\gg$ electroweak scale
$k_{\chi}^{2}$ $\ll 10^{-11}M_{p}^{2}$ $\ll\xi M^{2}$ (efficient preheating at the end of inflation)
Table 3: The viable values of the parameters in our model.